OpenStudy (anonymous):

Find the values of the six trigonometric functions of θ. Function Value: sin θ = 0 Constraint: π/2 ≤ θ ≤ 3π/2

OpenStudy (anonymous):

HELP ME PLEASE.

OpenStudy (anonymous):

you use your unit circle to find an angle whose sin is 0 the constraint tells you that the angle(s) is located in quadrant 2 and/or 3 So, at 180 degrees, or pi, the coordinates are (-1, 0) so sin is 0

OpenStudy (anonymous):

how do u find the six trigonometric functions using that though?

OpenStudy (cwrw238):



OpenStudy (cwrw238):

the angle is pi radians (180 degrees) cos pi = 1 tan pi = 0 csn u find the other 3?

OpenStudy (cwrw238):

recall that cot = 1/tan, sec = 1/cos and csc = 1 / sin

OpenStudy (cwrw238):